OpenStudy (anonymous):

Find the average value of f(x) = sin^2x+ cos^2x+x on the interval [0,pi]

hartnn (hartnn):

Average value over interval (a,b) is just \(\Large f_{avg} = \dfrac{1}{b-a}\int_a^b f(x)dx\)

OpenStudy (anonymous):

i broke it into 3 integrands and solved got pi/2,pi/2, and pi^2/2

OpenStudy (anonymous):

and then yes put it over pi but according to wolfram the solution is (pi+2)/2

hartnn (hartnn):

why not just use the fact that \(\Large \sin^2x+ \cos^2x =1\)

hartnn (hartnn):

your f(x) is just 1+x

OpenStudy (anonymous):

thanks for making me feel like a jack retricelol. i should really study those trig identities

hartnn (hartnn):

lol, you're learning, it happens :)

OpenStudy (anonymous):

ok still somewhat incorrect. that still gives pi+pi/2

OpenStudy (anonymous):

over pi for the avg

hartnn (hartnn):

integral of x is x^2/2 so you mean pi+pi^2/2 ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wolfram has 1/2pi(pi+2)

OpenStudy (anonymous):

or (2+pi)

hartnn (hartnn):

1+pi/2 wolfram and you, both get this

hartnn (hartnn):

pi+pi^2/2 this divided by pi is 1+pi/2 or (2+pi)/2

hartnn (hartnn):

i don't see any mismatch

OpenStudy (anonymous):

no your absolutely correct

hartnn (hartnn):

ask if anymore doubts :)

OpenStudy (anonymous):

40 gold stars for you today. snaps for hartnn

hartnn (hartnn):

lol, thank you :D